OpenStudy (anonymous):

Wilma bought 4 boxes of Crunch-A-Lot cereal. One out of every 5 boxes has a coupon for a free box of Crunch-A-Lot. What is the probability that Wilma got at least 2 coupons?

OpenStudy (anonymous):

4/5 chance of getting thre coupon

OpenStudy (anonymous):

The answer is about 2.6% but i don't know how they got that answer..

OpenStudy (anonymous):

let me do it again

OpenStudy (anonymous):

coupon in 1 box = 1/5 no coupon in 1 box = 4/5 no coupon in 5 boxes = (no coupon in 1 box)^5 no coupon in 4 boxes = (no coupon in 1 box)^4 exactly 1 coupon in 5 boxes = no coupon in 4 boxes * coupon in 1 box neither no coupon in 5 boxes nor exactly 1 coupon in 5 boxes = 1 - no coupon in 5 boxes - exactly 1 coupon in 5 boxes

OpenStudy (anonymous):

= 1 - (4/5)^5 - (4/5)^4 * (1/5)

OpenStudy (anonymous):

looks good to me at least 2 means not none and not one

OpenStudy (anonymous):

The answer is about 2.6% but i don't know how they got that answer. coupon in 1 box = 1/5 no coupon in 1 box = 4/5 no coupon in 5 boxes = (no coupon in 1 box)^5 no coupon in 4 boxes = (no coupon in 1 box)^4 exactly 1 coupon in 5 boxes = no coupon in 4 boxes * coupon in 1 box neither no coupon in 5 boxes nor exactly 1 coupon in 5 boxes = 1 - no coupon in 5 boxes - exactly 1 coupon in 5 boxes = 1 - (4/5)^5 - (4/5)^4 * (1/5) .

OpenStudy (anonymous):

wait, I solved it for if she bought 5 boxes, and forgot to account for the fact that there's more than one way for her to win only one 1 - (4/5)^4 - 4*(4/5)^3 * (1/5) =18.08% 2.6% is the right answer to "What's the probability she got 3 coupons?"